# CleanAgent: Automating Data Standardization with LLM-based Agents

### v1

@@ Abstract.
Data standardization is a crucial part in data science life cycle. While tools like Pandas offer robust functionalities, their complexity and the manual effort required for customizing code to diverse column types pose significant challenges. Although large language models (LLMs) like ChatGPT have shown promise in automating this process through natural language understanding and code generation, it still demands expert-level programming knowledge and continuous interaction for prompt refinement. To solve these challenges, our key idea is to propose a Python library with declarative, unified APIs for standardizing column types, simplifying the LLM’s code generation with concise API calls. We first propose Dataprep.Clean which is written as a component of the Dataprep Library, offers a significant reduction in complexity by enabling the standardization of specific column types with a single line of code. Then we introduce the CleanAgent framework integrating Dataprep.Clean and LLM-based agents to automate the data standardization process. With CleanAgent , data scientists need only provide their requirements once, allowing for a hands-free, automatic standardization process. To demonstrate the practical utility of CleanAgent , it has been developed as a user-friendly web application, allowing VLDB attendees to interact with it using real-world datasets.
@@ 1. Introduction
Example 1. We illustrate the data standardization task in Figure 1. Given an input table which needs to be standardized, it is obvious that data in the “Admission Date” column and the “Country of Birth” column are with different formats. Data in the three cells of “Admission Date” column includes three different date formats. “2021-01-01” follows the “YYYY-MM-DD” format, “2021.01.01” follows the “YYYY.MM.DD” format and “20/09/01” follows the “YY/MM/DD” format, where “YYYY / YY” represents year, “MM” represents month, “DD” represents day. The goal of data standardization is to unify the format of data in each column in meeting data scientists’ requirements and to get the standardized table . In Figure 1, the data scientists input their requirement to standardize “Admission Date” with the “MM/DD/YYYY” format. And in , data in the three cells of “Admission Date” column follows only one date format, i.e. the “MM/DD/YYYY” format.
@@
Recently, the emergence of large language models (LLMs), especially ChatGPT, has shown potential in revolutionizing this process. By leveraging their natural language understanding and code generation ability, these models could significantly aid data scientists by autonomously generating standardization code in response to conversational prompts. However, this method still necessitates detailed prompt crafting and often involves multi-turn dialogues, which limits efficiency and practicality.
To overcome these limitations, our key idea is to introduce a Python library involving declarative and unified APIs specifically designed for standardizing different column types. This idea simplifies the LLM’s task to converting natural language (NL) instructions into succinct, declarative API calls. Such an approach streamlines the LLM’s code generation process, requiring just a few lines of code,
@@
To solve C1, we propose the type-specific Clean module in Dataprep Library111https://github.com/sfu-db/dataprep, named Dataprep.Clean222https://github.com/sfu-db/dataprep/tree/develop/dataprep/clean. By observing the common steps of data standardization for specific column types, we design unified APIs clean_type(df, column_name, target_format), where the type represents the desired standardization type, such as date, address and phone, etc. These unified APIs offer enhanced expressiveness compared to raw Pandas code, reducing the complexity of standardizing specific column types. With Dataprep.Clean, data scientists can standardize one column type with only one line of code.
@@
Example 2. Continuing with Example 1. Given an input table which needs to be standardized and the data scientists’ requirements, the CleanAgent first recognizes that the “Admission Date” column belongs to the date type, and the “Country of Birth” column belongs to the country type. According to the column-type annotation results, the CleanAgent generates and executes Python code for standardization by calling “clean_date” and “clean_type” functions, then returns the standardized table .
CleanAgent is built as a web application. We allow the VLDB attendees to choose sample data and communicate with CleanAgent for standardization. We provide the demonstration video which can be found on Youtube333https://youtu.be/fSYXVM6qeqM.
To summarize, we make the following contributions: (1) We propose Dataprep.Clean, an open-sourced library showing reducing the complexity of implementing standardization for specific column types with type-specific standardization functions. (2) We propose CleanAgent , which automates the data standardization process by combining both the advantages of Dataprep.Clean and LLM-based Agents. (3) We deploy CleanAgent as a web application with user-friendly interface and demonstrate its utility. We also open-sourced the implementation of CleanAgent on GitHub444https://github.com/sfu-db/CleanAgent.
@@ 4. Demonstration Scenarios
5. Conclusion and Future Work
In this demo paper, we proposed CleanAgent to automate the data standardization process with Dataprep.Clean and LLM-based Agents. We implemented CleanAgent as a web service to visualize the conversations among agents. Other tasks in data science life cycle such as data cleaning and data visualization can also be completed by LLM-based agents (Xue et al., 2023). In the future, it is promising that the data science life cycle can be automatically planned and completed by LLM-based agents’ cooperation.

### v2

@@ Abstract.
Data standardization is a crucial part in data science life cycle. While tools like Pandas offer robust functionalities, their complexity and the manual effort required for customizing code to diverse column types pose significant challenges. Although large language models (LLMs) like ChatGPT have shown promise in automating this process through natural language understanding and code generation, it still demands expert-level programming knowledge and continuous interaction for prompt refinement. To solve these challenges, our key idea is to propose a Python library with declarative, unified APIs for standardizing different column types, simplifying the LLM’s code generation with concise API calls. We first propose Dataprep.Clean which is written as a component of the Dataprep Library, offers a significant reduction in complexity by enabling the standardization of specific column types with a single line of code. Then we introduce the CleanAgent framework integrating Dataprep.Clean and LLM-based agents to automate the data standardization process. With CleanAgent , data scientists need only provide their requirements once, allowing for a hands-free, automatic standardization process. To demonstrate the practical utility of CleanAgent , it has been developed as a user-friendly web application, allowing VLDB attendees to interact with it using real-world datasets.
@@ 1. Introduction
Example 1. We illustrate the data standardization task in Figure 1. Given an input table which needs to be standardized, it is obvious that data in the “Admission Date” column and the “Address” column are with different formats. Data in the three cells of “Admission Date” column includes two different date formats. The goal of data standardization is to unify the format of data in each column in meeting data scientists’ requirements and to get the standardized table . In Figure 1, the data scientists input their requirement to standardize “Admission Date” with the “MM/DD/YYYY HH:MM:SS” format. And in , data in the three cells of “Admission Date” column follows only one date format, i.e. the “MM/DD/YYYY HH:MM:SS” format.
@@
Example 2. Still considering the data standardization task in Figure 1. For standardizing “Admission Date” and “Address” , data scientists need to write the datetime standardization code for “Admission Date” and address standardization code for “Address” by using regex. The example standardization code for “Address” is shown as follows.
If the input table has other column types such as email and IP addresses, data scientists also need to write standardization code specially for email type and IP address type, which is time-consuming.
Recently, the emergence of LLMs, especially ChatGPT, has shown potential in revolutionizing this process. By leveraging their natural language understanding and code generation ability, these models could significantly aid data scientists by autonomously generating standardization code in response to conversational prompts. However, this method still necessitates detailed prompt crafting and often involves multi-turn dialogues (Chen et al., 2023) for standardizing column types in one by one, which limits efficiency and practicality.
To overcome these limitations, our key idea is to introduce a Python library involving declarative and unified APIs specifically designed for standardizing different column types. This idea simplifies the LLM’s task to converting natural language (NL) instructions into succinct, declarative API calls. Such an approach simplifies the LLM’s code generation process for data standardization, requiring just a few lines of code.
@@
To solve C1, we propose the type-specific Clean module in Dataprep Library, named Dataprep.Clean. By observing the common steps of data standardization for specific column types, we design unified APIs clean_type(df, column_name, target_format), where the type represents the desired standardization type, such as date, address and phone, etc. These unified APIs offer enhanced expressiveness compared to raw Pandas code, reducing the complexity of standardizing specific column types. With Dataprep.Clean, data scientists can standardize one column type with only one line of code.
@@
Example 3. Continuing with Example 1. Given an input table which needs to be standardized and the data scientists’ requirements, the CleanAgent first recognizes that the “Admission Date” column belongs to the date type, and the “Address” column belongs to the address type. According to the column-type annotation results, the CleanAgent generates and executes Python code for standardization by calling “clean_date” and “clean_address” functions, then returns the standardized table .
CleanAgent is built as a web application. We allow the VLDB attendees to choose sample data and communicate with CleanAgent for standardization. We provide the demonstration video which can be found on Youtube.
To summarize, we make the following contributions: (1) We propose Dataprep.Clean, an open-sourced library showing reducing the complexity of implementing standardization for specific column types with type-specific standardization functions. (2) We propose CleanAgent , which automates the data standardization process by combining both the advantages of Dataprep.Clean and LLM-based Agents. (3) We deploy CleanAgent as a web application with user-friendly interface and demonstrate its utility. We also open-sourced the implementation of CleanAgent on Github.
@@ 4. Demonstration Scenarios
5. Conclusion
In this demo paper, we proposed CleanAgent to automate the data standardization process with Dataprep.Clean and LLM-based Agents. We implemented CleanAgent as a web service to visualize the conversations among agents. Other tasks in data science life cycle such as data cleaning and data visualization can also be completed by LLM-based agents (Xue et al., 2023).